OpenStudy (anonymous):

a. Write the value(s) of the variable that make a denominator zero. b. Solve the equation 2/x-2 = x/x-2 -2

OpenStudy (anonymous):

is this \[\frac{2}{x-2}=\frac{x}{x-2}-2\]?

OpenStudy (anonymous):

if so the answer to question a) is x = 2

OpenStudy (anonymous):

Yes. and ok how would you get that?

OpenStudy (anonymous):

set \[x-2=0\] get \[x=2\]